OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (mathmate):

Can you find the slope of the new line?

OpenStudy (mathmate):

Better still, can you first convert the equation to standard slope intercept form?

OpenStudy (anonymous):

i still dont understand can you please help me

OpenStudy (mathmate):

Convert the equation to slope-intercept form means: from 2x − 3y = 3 we write y=(2x-3)/3 or y=(2/3)x-1 this will show clearly the original equation has a slope of (2/3). Can you tell me the slope of a line which is perpendicular to the ine y=(2/3)x-1

OpenStudy (anonymous):

well thank you , i understand

OpenStudy (mathmate):

You're welcome!